OpenStudy (anonymous):

if line A is parallel to 3x+4y=6 and passes through (3,5) then what is the equation of line A

OpenStudy (anonymous):

Can you put that in slope intercept form y=mx+b?

OpenStudy (anonymous):

will it be y= -3/4 x +6

OpenStudy (anonymous):

Close, you have to divide the 6 by 4 also. y=-3/4x+3/2 can you now see what the slope is?

OpenStudy (anonymous):

-3/4

OpenStudy (anonymous):

Yep, and parallel lines have the same slope. So use y-y1=m(x-x1) to get your equation. (x1,y1) is the point it have you and m is the slope, just plug those in and simplify

OpenStudy (anonymous):

thanx